OpenStudy (anonymous):

evaluate the logarithm. log5 (1/25)

OpenStudy (kainui):

@Jhannybean

OpenStudy (jhannybean):

\[\large \log_a (x) = \frac{\log_b (x)}{\log_b (a)}\]

OpenStudy (jhannybean):

@Kainui

OpenStudy (anonymous):

(1/25)/5????

OpenStudy (jhannybean):

close, where are your function denotations?

OpenStudy (anonymous):

what do you mean?

OpenStudy (kainui):

Well I wouldn't use that, I'd use the fact that: \[\log_5(\frac{1}{25}) = \log_5(5^{-2})\] and so log base 5 asks us what exponent is on 5? that exponent is -2, so the answer is -2.

OpenStudy (jhannybean):

\[\log_5 \left(\frac{1}{25}\right) = \frac{\log\left(\frac{1}{25}\right)}{\log(5)} = \frac{\log(1) -\log(5^2)}{\log(5)}=...\]